OpenStudy (anonymous):

(4x+8)/(2x) - (x)/(x-4) = Once you find the answer discuss why the degree of the resulting denominator did not change from your expression’s degree. PLEASE PLEASE PLEASE HELP!!!!!!!!!!!

OpenStudy (imstuck):

Well the way you solve this is to factor the numerator. Like this:

OpenStudy (imstuck):

When you do this you will get 4(x + 2)/2x. Looking at it more clearly,

OpenStudy (neer2890):

\[\frac{ 4x+8 }{ 2x }-\frac{ x }{ x-4 }=\frac{ (4x+8)(x-4)-x(2x) }{ 2x(x-4) }\]

OpenStudy (imstuck):

You can cancel between the 2 and the 4; the 2 from the 2x, not the one inside the parenthesis. That leaves an x only in the denominator. That is why the denominator's resulting degree did not change. You CANNOT cancel what is inside the parenthesis. You can' touch it! Only what is outside is legal!

OpenStudy (imstuck):

Oh my...the problem is not just what I typed...it goes on and on...I didn't see that part as did neer2890! Oops!

OpenStudy (neer2890):

\[=\frac{ 4x ^{2}-16x+8x-32-2x ^{2} }{ 2x(x-4) }=\frac{ 4x ^{2}-2x ^{2}-16x+8x-32 }{ 2x(x-4) }\]

OpenStudy (anonymous):

So @neer2890 is correct?

OpenStudy (anonymous):

Is that the final simplified answer

OpenStudy (neer2890):

sorry.. i think i got a mistake let me correct it.

OpenStudy (anonymous):

ok

OpenStudy (neer2890):

\[=\frac{ 2x ^{2}-8x-32 }{ 2x(x-4) }=\frac{ 2(x ^{2}-4x-16 )}{ 2x(x-4) }=\frac{ x ^{2}-4x-16 }{ x(x-4) }\]

OpenStudy (anonymous):

Okay Thank you

OpenStudy (neer2890):

you're welcome..:) do you know the 2nd part?

OpenStudy (anonymous):

No i don't. Would you be able to help me? :)

OpenStudy (neer2890):

ok. you can see that the numerator of the fraction is a quadratic function of the form\[ax ^{2}+bx+c\] where a=1 ,b=-4 and c=-16

OpenStudy (anonymous):

Yes i see

OpenStudy (neer2890):

do you know how to factorize the quadratic functions?

OpenStudy (anonymous):

Um...not really

OpenStudy (neer2890):

to factorize quadratic functions of the form \[ax ^{2}+bx+c\] we first multiply first and 3rd term which gives us \[acx ^{2}\] and now we have to factor this \[acx ^{2}\] into two parts such that it their addition becomes bx and their multiplication becomes \[acx ^{2}\]

OpenStudy (anonymous):

Are you talking about how the resulting denominator did not change from the expression degree? the to figure out the answer to that question

OpenStudy (neer2890):

yes

OpenStudy (anonymous):

What exactly is that question asking of me? i really dont understand at all

OpenStudy (neer2890):

ok. let me complete then you will understand what's going on.

OpenStudy (anonymous):

okay

OpenStudy (neer2890):

for e.g., \[x ^{2}-4x-32\] first multiply x^2 and -32 which gives us\[-32x ^{2}\] now we have to factor this \[-32x ^{2}\] into two parts such that their addition becomes -4x and their multiplication becomes \[-32x ^{2}\]

OpenStudy (neer2890):

so +4x and -8x are the factors of this function. and we can write it as \[x ^{2}-4x-32\] =\[x ^{2}+4x-8x-32\] =x(x+4)-8(x+4) =(x+4)(x-8)

OpenStudy (neer2890):

this is called factorization of a quadratic polynomial but you can see that we can not factorize \[x ^{2}-4x-16\]

OpenStudy (anonymous):

So theresulting denominator did not change from your expressions degree because you cannot factorize?

OpenStudy (anonymous):

???

OpenStudy (neer2890):

if we can't factorize the numerator then how come the denominator changes. because if we can factor the numerator and one of the factor is (x-4) then it cancels out and the denominator changes. But here, we can't factor the numerator that's why the denominator remains same.